OpenStudy (anonymous):

What is the area of a sector with radius 10" and measure of arc equal to 45°? help me please

OpenStudy (anonymous):

see circle area is pir^2 which is pix100....when theta = 90 its quarter pi25 is the area but 45 is half so the area must be 25/2*pi

OpenStudy (anonymous):

Sorry, writing the final answer: \[45°=\frac{\pi}{4}rad\] \[A=\frac{\theta \times r^2}{2}\] With theta in radians \[A=\frac{\pi \times 100}{2 \times 4}=\frac{25\times \pi}{2}\] in square inches

hero (hero):

\[\frac{x}{{\pi}r^2} = \frac{45}{360}\] \[\frac{x}{{\pi}10^2} = \frac{45}{360}\]

hero (hero):

Cross Multiply, solve for x

hero (hero):

\[360x = 4500\pi \] \[x = \frac{4500{\pi}} {360} \]

hero (hero):

x = 39.2 units squared

hero (hero):

or x = 25pi/2 units squared